# Quadratic Time–Frequency Analysis of Vibration Signals for Diagnosing Bearing Faults

### v1

@@ Abstract
Index Terms:
Bearing fault, damage detection, deep learning, time-frequency analysis, variable speed.
@@ II-A Signal model
where $\boldsymbol{h}(t)$ holds the bearing normal vibration (healthy patterns), $\boldsymbol{f}(t)$ is the fault-specific waveform, $A$ is the fault amplitude, $f_{c}$ is the bearing resonance frequency, $\beta$ is the damping characteristic factor, *$\underset{(t)}{*}$ denotes convolution in time, $\delta(t)$ is the Dirac function, $f_{\text{fault}}(t)$ denotes the time-varying fault characteristic frequency, and $K$ is the number of fault impulses, or knocks, within the signal duration $T$. The characteristic frequencies that appear in the vibration signal of a faulty bearing, assuming a stationary outer race and a rotating inner race, are mainly the ball pass frequency of the inner race, ball pass frequency of the outer race, and ball spinning frequency, which are expressed by [43]:
@@
Figure 2: Example synthetic bearing vibration signal with an inner race fault following Eq. (4). (a) shows the linearly changing rotational speed in red and its pulse frequency modulation (PFM) encoding in black, (b) depicts the generated vibration signal, (c) illustrates its power spectrum, and (d)-(f) show the compact kernel distribution (CKD) high accuracy, compared to the Wigner-Ville distribution (WVD), and its high resolution, compared to the Spectrogram, when representing the vibration signal in the time-frequency domain. The signal model parameters are set to $f_{s}=1$ kHz, $T=5$ seconds, $A=1$, $f_{c}=250$ Hz, $\beta=0.1$, $n=9$, $D_{r}=7.9$ mm, $D_{p}=38.5$ mm, $\theta=0\degree$, and $f_{r}(t)$ is a triangular wave rising from 100 rpm to 500 rpm and back to 100 rpm. The CKD parameters are set to $c=1$, $D=0.1$, $E=0.1$, and the Spectrogram is computed at full resolution using a 0.3 seconds Hamming window. The TFRs are plotted on a logarithmic scale to ease visualizing the fault-induced waveforms around the bearing’s resonance at $250$ Hz (-3 dB to 2.2 dB).
@@ II-B Quadratic time-frequency representation
where $\boldsymbol{\rho}_{\boldsymbol{z}}(t,f)$ holds smoothed TFDs and **$\underset{(t,f)}{**}$ denotes the convolution in time and frequency. This expression can be written in a more convenient space, called the Doppler-lag domain, where convolutions become multiplications [18], i.e.:
@@ III-A Evaluation of bearing fault diagnosis
Table I illustrates the testing performance of the TF-CNN model and compares it to the two competing methods (T-CNN [4] and PIResNet [47]) across different SNR levels. The results reveal a consistent superiority of the proposed model over the other techniques across all metrics and SNR levels. Additionally, they show that the boost in performance becomes more pronounced when the level of noise increases in the vibration signal. For instance, when utilizing clean data, the TF-CNN outruns PIResNet and T-CNN with 2.2% and 1.6% point gains in accuracy, respectively. Nonetheless, these gains increase to approximately 19% and 15% when dealing with severely degraded vibration signals at -5 dB. This anticipated improvement in performance can be attributed to the uniform distribution of noise within the TF domain. Consequently, quadratic TFRs are well-suited to unveil distinctive patterns associated with the different bearing faults while concurrently mitigating the influence of noise.
@@ III-B The t-SNE and Grad-CAM analysis
A two-dimensional representation of the TF-CNN learned features is illustrated in Fig. 6 using t-SNE. The results validate the model’s ability to learn state-relevant features in case of clean or noisy vibration data. Specifically, Figs. 6a and 6b show clear inter-class separation among all classes while Fig. 6c depicts a slight overlap between the Normal and Outer classes. Nonetheless, this overlap escalates when heavy noise is introduced as shown in Fig. 6d. Moreover, the t-SNE results reveal that the Normal and Outer states are the most alike when degraded with heavy noise, while the Inner and Ball classes are the most distinct from the rest. These observations are aligned with the findings in Fig. 5 and justify the slight performance imbalance at -5 dB. We hypothesize that this alikeness is linked to the characteristic frequency of outer race faults being the lowest among the faults observed. This particular frequency characteristic causes their harmonics to sway less from the resonance of the bearing. Consequently, in the TF domain, outer race faults tend to overlap more with the normal state, and substantial noise can significantly obscure their distinguishing marks.
@@
Furthermore, Fig. 7 illustrates the Grad-CAM outcomes for all the classes along with the rotational speeds for the first 8 seconds of clean vibrations. The visualizations validate our claim that the diagnosis of bearing faults is beyond the scope of basic harmonic analysis. In particular, they underscore the necessity for dynamic features derived from the joint TF domain and indicate apparent limitations when using time-only or frequency-only representations. Additionally, the results showcase the versatility of the proposed model in acquiring fault-relevant non-stationary features, i.e., important descriptors are extracted from TF regions that evolve in time and frequency; see Figs. 7b-7d. Finally, by associating the rotational speed profiles with the highlighted TF regions, one notes a moderate correspondence between the region’s frequency support and the motor’s speed, i.e., the regions are temporally coupled with the rotational speeds. However, the outcomes also demonstrate a need for improving the proposed model’s attention to generate predictions with higher correlations to the operating variable speeds.

### v3

# Fusion of Quadratic Time–Frequency Analysis and Convolutional Neural Networks to Diagnose Bearing Faults Under Time-Varying Speeds

@@ II-A Signal model
where $\boldsymbol{h}(t)$ holds the bearing normal vibration (healthy patterns), $\boldsymbol{f}(t)$ is the fault-specific waveform, $A$ is the fault amplitude, $f_{c}$ is the bearing resonance frequency, $\beta$ is the damping characteristic factor, $\underset{(t)}{*}$ denotes convolution in time, $\delta(t)$ is the Dirac function, $f_{\text{fault}}(t)$ denotes the time-varying fault characteristic frequency, and $K$ is the number of fault impulses, or knocks, within the signal duration $T$. The characteristic frequencies that appear in the vibration signal of a faulty bearing, assuming a stationary outer race and a rotating inner race, are mainly the ball pass frequency of the inner race, ball pass frequency of the outer race, and ball spinning frequency, which are expressed by [43]:
@@
Figure 2: Example synthetic bearing vibration signal with an inner race fault following Eq. (4). (a) shows the linearly changing rotational speed in red and its pulse frequency modulation (PFM) encoding in black, (b) depicts the generated vibration signal, (c) illustrates its power spectrum, and (d)-(f) show the compact kernel distribution (CKD) high accuracy, compared to the Wigner-Ville distribution (WVD), and its high resolution, compared to the Spectrogram, when representing the vibration signal in the time-frequency domain. The signal model parameters are set to $f_{s}=1$ kHz, $T=5$ seconds, $A=1$, $f_{c}=250$ Hz, $\beta=0.1$, $n=9$, $D_{r}=7.9$ $m$, $D_{p}=38.5$ $m$, $\theta=0\degree$, and $f_{r}(t)$ is a triangular wave rising from 100 rpm to 500 rpm and back to 100 rpm. The CKD parameters are set to $c=1$, $D=0.1$, $E=0.1$, and the Spectrogram is computed at full resolution using a 0.3 seconds Hamming window. The TFRs are plotted on a logarithmic scale to ease visualizing the fault-induced waveforms around the bearing’s resonance at $250$ Hz (-3 dB to 2.2 dB).
@@ II-B Quadratic time-frequency representation
where $\boldsymbol{\rho}_{\boldsymbol{z}}(t,f)$ holds smoothed TFDs and $\underset{(t,f)}{**}$ denotes the convolution in time and frequency. This expression can be written in a more convenient space, called the Doppler-lag domain, where convolutions become multiplications [18], i.e.:
@@ III-A Evaluation of bearing fault diagnosis
Table I illustrates the testing performance of the TF-CNN model and compares it to the two competing methods (T-CNN [4] and PIResNet [47]) across different SNR levels. The results reveal a consistent superiority of the proposed model over the other techniques across all metrics and SNR levels. Additionally, they show that the boost in performance becomes more pronounced when the level of noise increases in the vibration signal. For instance, when utilizing clean data, the TF-CNN outruns PIResNet and T-CNN with 2.2% and 1.6% point gains in accuracy, respectively. Nonetheless, these gains increase to approximately 19% and 15% when dealing with severely degraded vibration signals at -5 dB. This anticipated improvement in performance can be attributed to the uniform distribution of noise within the TF lattice. Consequently, quadratic TFRs are well-suited to unveil distinctive patterns associated with the different bearing faults while concurrently mitigating the influence of noise.
@@ III-B The t-SNE and Grad-CAM analysis
A two-dimensional representation of the TF-CNN learned features is illustrated in Fig. 6 using t-SNE.
@@
The results validate the model’s ability to learn state-relevant features in case of clean or noisy vibration data. Specifically, Figs. 6a and 6b show clear inter-class separation among all classes while Fig. 6c depicts a slight overlap between the Normal and Outer classes. Nonetheless, this overlap escalates when heavy noise is introduced as shown in Fig. 6d. Moreover, the t-SNE results reveal that the Normal and Outer states are the most alike when degraded with heavy noise, while the Inner and Ball classes are the most distinct from the rest. These observations are aligned with the findings in Fig. 5 and justify the slight performance imbalance at -5 dB. We hypothesize that this alikeness is linked to the characteristic frequency of outer race faults being the lowest among the faults observed. This particular frequency characteristic causes their harmonics to sway less from the resonance of the bearing. Consequently, in the TF domain, outer race faults tend to overlap more with the normal state, and substantial noise can significantly obscure their distinguishing marks.
Furthermore, Fig. 7 illustrates the Grad-CAM outcomes for all the classes along with the rotational speeds for the first 8 seconds of clean vibrations. The visualizations validate our claim that the diagnosis of bearing faults is beyond the scope of basic harmonic analysis. In particular, they underscore the necessity for dynamic features derived from the joint TF domain and indicate apparent limitations when using time-only or frequency-only representations. Additionally, the results showcase the versatility of the proposed model in acquiring fault-relevant non-stationary features, i.e., important descriptors are extracted from regions that evolve in time and frequency; see Figs. 7b-7d. Finally, by associating the rotational speed profiles with the highlighted TF regions, one notes a moderate correspondence between the region’s frequency support and the motor’s speed, i.e., the regions are temporally coupled with the rotational speeds. However, the outcomes also demonstrate a need for improving the proposed model’s attention to generate predictions with higher correlations to the operating variable speeds.
@@
III-C The misclassifications link to speed
The connection between the TF-CNN false decisions at 0 dB SNR and the motor rotational speed is explored in Fig. 8. First, the temporal locality of the model’s misclassifications is presented in Fig. 8a for all the classes. The locale is overlaid on the rotational speed profile for the entire 35 minutes of data, i.e., all the testing data from all the folds. Moreover, Fig. 8b shows the motor speed probability distribution given false predictions and compares it to the speed’s unconditional distribution (baseline) for all the bearing health states. Interestingly, the results reveal that misclassifications are most likely to occur at speeds below 1,200 RPM (the lower 1/3 of the entire speed range) where 70%, 100%, 69%, and 73% of all misclassifications exist for the Normal, Inner, Outer, and Ball classes, respectively. Additionally, the conditional distributions show that the median speed for which the TF-CNN fails to predict the bearing health state correctly is 1,046 for Normal, 840 for Inner, 974 for Outer, and 992 for the Ball state.
The observed skewness towards low speeds can be attributed to the spectral signature of fault-induced vibrations being very close to the bearing’s resonance. In other words, the frequency of fault-induced harmonics varies less at those speeds; hence, their TFRs are more likely to be misidentified. Moreover, the constructive accumulation of fault-induced impulses, as described in Eq. (4), is linearly dependent on the rotational speed, i.e., the envelope of the fault signal co-varies with the speed; see Fig. 2b for illustration. Consequently, the ubiquity of the fault signal becomes obscure at severe noise conditions and low operational speeds. These conjectures are aligned with the t-SNE analysis in Fig. 6 and anticipate an increase in the fault-conditioned median speed at lower SNR levels. We verified this prediction at -5 dB SNR and found that the median speed is 1,252 for Normal (20% increase), 900 for Inner (7% increase), 1,314 for Outer (35% increase), and 1,244 for the Ball state (25% increase).
(a) The misclassification temporal locale w.r.t. speed.
(b) The probability distribution of rotational speeds given misclassification.
Figure 8: The link between misclassification, at 0 dB SNR, and rotational speed for the different bearing health states. (a) shows the temporal locale of misclassifications in relation to the motor’s speed profile, and (b) compares the speed conditional probability distributions given false decisions to the baseline, i.e., the unconditional probability distribution of the time-varying speed. The baseline speed profiles in (a) are interpolated by the modified Akima cubic Hermite interpolation method to match the decisions’ sampling time. The probability distributions in (b) are smoothed using normal kernels and plotted along with their 25% (Q1), 50% (Q2), and 75% (Q3) percentiles.